# Towards a Functionally Complete and Parameterizable TFHE Processor

### v1

@@ 5.2. Baseline for Comparisons
While prior works employ slightly different Boolean parameter sets, typically $n=500$, $N=1024$, $k=1$, $\ell=2$, $\log\beta=10$, and $q=2^{32}$, ALT adopts modified parameters with $\ell=1$, $\log\beta=16$, and a composite modulus. To enable fair comparison, we normalize the decryption failure rate (DFR) by assuming that repeating a computation squares the DFR. Accordingly, we scale it by $z=\frac{\log_{2}(\text{DFR}_{x})}{-64}$ to match the standard DFR of $2^{-64}$ (Zama, 2022a). This normalization inherently compensates for parameter variations that affect the DFR.
@@
HLW (Hartshorn et al., 2020) 64 179 1294 7.23 138 9 K 9 K 128 - 8
KL (Kong and Li, 2024) 52 433 697 1.61 621 33 K 44 K 160 - 40
YYKKP (Ye et al., 2021) 32 215 198 0.92 1086 94 K 105 K 640 - 80
Proteus (Hirner et al., 2024) 32 150 1100 7.3 300 7 K 3 K 36 - 2
CM (Mert et al., 2020) 28 125 490 3.90 255 16 K 14 K 56 - 24
@@ 5.4. NTT Performance Comparison
In order to illustrate the efficiency of our implementation, we compare our 64-bit negacyclic NTT with prior FPGA-based designs, summarized in Table 3. Among them, YYKKP (Ye et al., 2022) and KL (Kong and Li, 2024) denote the NTT implementations of the respective TFHE accelerators, while Proteus (Hirner et al., 2024), HLW (Hartshorn et al., 2020), and CM (Mert et al., 2020) provide standalone 1024-point NTTs. All use Virtex-family FPGAs with comparable resource architectures. Our design, operating at 600 MHz with $\mathcal{T}=2$, achieves the highest frequency, throughput (NTTs/ms) and lowest latency among all works except for YYKKP, which attains slightly lower latency but requires over $7\times$ more DSPs, $5\times$ more FFs, and $7\times$ more LUTs while using half the bit width. Compared to HLW, which uses the same modulus $p=2^{64}-2^{32}+1$, our design delivers $\sim 9\times$ higher throughput and $6\times$ lower latency. It also outperforms KL on the same FPGA and exceeds Proteus and CM by around $4\times$ in performance while maintaining competitive resource efficiency. Overall, our NTT demonstrates superior performance with resource balance and forms a high-speed backbone for polynomial multiplications within the blind rotation.
@@ 5.5. Full System Performance Comparison
Evaluation on Standard Parameter Set: Table 4 summarizes the results in terms of decryption failure rate (DFR), number of coefficients processed per clock cycle ($\mathcal{T}$), operating frequency, throughput in PBS/s (for 64-bit, 32-bit, and 32-bit normalized to DFR $2^{-64}$), latency, hardware resource utilization, and power consumption. For the sake of completeness, CPU and GPU implementations are also included as reference baselines in the comparison. Note that we take the DFR for ALT and FPT as reported by ALT (Hu et al., 2024).
Table 4. Comparison with literature results on the parameter set: $n=500$, $N=1024$, $k=1$, $\ell=2$ and $\log\beta=10$. * denotes estimated data. The comparison CPU is an AMD 7950X3D running tfhe-rs (Zama, 2022b). The comparison GPU 1 is an Nvidia L40 running tfhe-rs using the classic settings. The comparison GPU 2 is an RTX 3090 running cuFHE (Dai and Sunar, 2016), its results are from YKP’s benchmark (Ye et al., 2022). For the CPU and GPU 1 data, we took the Thermal Design Power (TDP) as their power consumption and averaged their performance over $10^{4}$ runs.
@@
In terms of energy efficiency, our processor achieves the second-best performance per watt, trailing only ALT. Regarding latency, our best configuration ranks behind ALT and FPT. Note that our 64-bit implementation is at a disadvantage since 32-bit multiplications require fewer cycles than our 64-bit multiplications. However, ALT attains the lowest latency primarily because of its smaller multiplications due to their composite NTT and additionally through employing bootstrapping key unrolling, which also reduces computation effort (Hu et al., 2024). Incorporating key unrolling in future versions could further improve our design. Our best performance-per-resource ratio is achieved at $\mathcal{T}=2$, while the highest absolute throughput occurs at $\mathcal{T}=32$. As $\mathcal{T}$ increases, frequency decreases due to longer interconnecting paths and higher routing congestion. Notably, a standalone NTT with $\mathcal{T}=8$ reaches nearly 600 MHz, indicating that inter-block routing rather than the NTT itself limits maximum frequency.
Table 5. Comparison with literature results on the parameter set: $n=800$, $N=16384$, $k=1$, $\ell=5$ and $\log\beta=6$ and DFR $2^{-64}$. * denotes estimated data. The comparison CPU is an AMD 7950X3D running TFHE-rs (Zama, 2022b) using all threads. The comparison GPU is an Nvidia L40 running TFHE-rs using the classic settings.
@@
Evaluation on Large Parameter Set: To demonstrate scalability and flexibility, we evaluate our TFHE processor using the demanding parameter set from Ohba et al. (Ohba et al., 2025), as shown in Table 5. Similar to previous experiments, we have included the CPU and GPU results for reference. Our design achieves over $30\times$ higher throughput and $10\times$ lower latency. Compared to a 16-core CPU (GPU) baseline, it delivers a $2.89\times$ ($1.16\times$) speedup on the smaller parameter set and $4.35\times$ ($1.12\times$) on the larger one, showing improved efficiency at scale against the CPU, and similar scalability to the L40 GPU.
In our architecture, parameter $k$ affects only buffer sizes, while $N$ primarily impacts BRAM and URAM blocks already partitioned into $\mathcal{T}$ sub-buffers. As long as these fit within FPGA RAM primitives, scaling parameters proportionally scale performance, such that they have a negligible performance impact. Technological differences also affect cross-platform comparisons: our FPGA is manufactured using a 20 nm process (AMD, 2025b), while comparison CPU and GPUs rely on 5–6 nm FinFET nodes (AMD, 2025a; techpowerup.com, 2025). Despite this disadvantage, our $\mathcal{T}=32$ design achieves over $5\times$ ($28\times$) lower latency and $2.8\times$ ($1.16\times$) higher throughput than a 16-core CPU (GPU) with similar or lower power consumption. In summary, these results highlight that our FPGA accelerator surpasses general-purpose processors in performance and efficiency, highlighting strong potential for ASIC-based TFHE solutions and practical privacy-preserving computation in real-world applications.

### v2

@@ 5.2. Baseline for Comparisons
While prior works employ slightly different Boolean parameter sets, typically $n=500$, $N=1024$, $k=1$, $\ell=2$, $\log\beta=10$, and $q=2^{32}$, ALT adopts modified parameters with $\ell=1$, $\log\beta=16$, and a composite modulus. To enable fair comparison, we normalize the decryption failure rate (DFR) by assuming that repeating a computation squares the DFR. Accordingly, we scale it by $z=\frac{\log_{2}(\text{DFR}_{x})}{-64}$ to match the standard DFR of $2^{-64}$ (concrete_library). This normalization inherently compensates for parameter variations that affect the DFR.
@@
HLW (MQT) 64 179 1294 7.23 138 9 K 9 K 128 - 8
KL (Tianqi_Kong) 52 433 697 1.61 621 33 K 44 K 160 - 40
YYKKP (Tian_Ye_NTT) 32 215 198 0.92 1086 94 K 105 K 640 - 80
Proteus (Proteus) 32 150 1100 7.3 300 7 K 3 K 36 - 2
CM (Ahmed_Can_Mert_2020) 28 125 490 3.90 255 16 K 14 K 56 - 24
@@ 5.4. NTT Performance Comparison
In order to illustrate the efficiency of our implementation, we compare our 64-bit negacyclic NTT with prior FPGA-based designs, summarized in Table 3. Among them, YYKKP (Tian_Ye) and KL (Tianqi_Kong) denote the NTT implementations of the respective TFHE accelerators, while Proteus (Proteus), HLW (MQT), and CM (Ahmed_Can_Mert_2020) provide standalone 1024-point NTTs. All use Virtex-family FPGAs with comparable resource architectures. Our design, operating at 600 MHz with $\mathcal{T}=2$, achieves the highest frequency, throughput (NTTs/ms) and lowest latency among all works except for YYKKP, which attains slightly lower latency but requires over $7\times$ more DSPs, $5\times$ more FFs, and $7\times$ more LUTs while using half the bit width. Compared to HLW, which uses the same modulus $p=2^{64}-2^{32}+1$, our design delivers $\sim 9\times$ higher throughput and $6\times$ lower latency. It also outperforms KL on the same FPGA and exceeds Proteus and CM by around $4\times$ in performance while maintaining competitive resource efficiency. Overall, our NTT demonstrates superior performance with resource balance and forms a high-speed backbone for polynomial multiplications within the blind rotation.
@@ 5.5. Full System Performance Comparison
Evaluation on Standard Parameter Set: Table 4 summarizes the results in terms of decryption failure rate (DFR), number of coefficients processed per clock cycle ($\mathcal{T}$), operating frequency, throughput in PBS/s (for 64-bit, 32-bit, and 32-bit normalized to DFR $2^{-64}$), latency, hardware resource utilization, and power consumption. For the sake of completeness, CPU and GPU implementations are also included as reference baselines in the comparison. Note that we take the DFR for ALT and FPT as reported by ALT (ALT).
Table 4. Comparison with literature results on the parameter set: $n=500$, $N=1024$, $k=1$, $\ell=2$ and $\log\beta=10$. * denotes estimated data. The comparison CPU is an AMD 7950X3D running tfhe-rs (tfhe_rs). The comparison GPU 1 is an Nvidia L40 running tfhe-rs using the classic settings. The comparison GPU 2 is an RTX 3090 running cuFHE (cuFHE), its results are from YKP’s benchmark (Tian_Ye). For the CPU and GPU 1 data, we took the Thermal Design Power (TDP) as their power consumption and averaged their performance over $10^{4}$ runs.
@@
In terms of energy efficiency, our processor achieves the second-best performance per watt, trailing only ALT. Regarding latency, our best configuration ranks behind ALT and FPT. Note that our 64-bit implementation is at a disadvantage since 32-bit multiplications require fewer cycles than our 64-bit multiplications. However, ALT attains the lowest latency primarily because of its smaller multiplications due to their composite NTT and additionally through employing bootstrapping key unrolling, which also reduces computation effort (ALT). Incorporating key unrolling in future versions could further improve our design. Our best performance-per-resource ratio is achieved at $\mathcal{T}=2$, while the highest absolute throughput occurs at $\mathcal{T}=32$. As $\mathcal{T}$ increases, frequency decreases due to longer interconnecting paths and higher routing congestion. Notably, a standalone NTT with $\mathcal{T}=8$ reaches nearly 600 MHz, indicating that inter-block routing rather than the NTT itself limits maximum frequency.
Table 5. Comparison with literature results on the parameter set: $n=800$, $N=16384$, $k=1$, $\ell=5$ and $\log\beta=6$ and DFR $2^{-64}$. * denotes estimated data. The comparison CPU is an AMD 7950X3D running TFHE-rs (tfhe_rs) using all threads. The comparison GPU is an Nvidia L40 running TFHE-rs using the classic settings.
@@
Evaluation on Large Parameter Set: To demonstrate scalability and flexibility, we evaluate our TFHE processor using the demanding parameter set from Ohba et al. (Ohba2025_nvme), as shown in Table 5. Similar to previous experiments, we have included the CPU and GPU results for reference. Our design achieves over $30\times$ higher throughput and $10\times$ lower latency. Compared to a 16-core CPU (GPU) baseline, it delivers a $2.89\times$ ($1.16\times$) speedup on the smaller parameter set and $4.35\times$ ($1.12\times$) on the larger one, showing improved efficiency at scale against the CPU, and similar scalability to the L40 GPU.
In our architecture, parameter $k$ affects only buffer sizes, while $N$ primarily impacts BRAM and URAM blocks already partitioned into $\mathcal{T}$ sub-buffers. As long as these fit within FPGA RAM primitives, scaling parameters proportionally scale performance, such that they have a negligible performance impact. Technological differences also affect cross-platform comparisons: our FPGA is manufactured using a 20 nm process (virtex_ultrascale_fpga_product_page), while comparison CPU and GPUs rely on 5–6 nm FinFET nodes (7950x3d_datasheet; L40_datasheet). Despite this disadvantage, our $\mathcal{T}=32$ design achieves over $5\times$ ($28\times$) lower latency and $2.8\times$ ($1.16\times$) higher throughput than a 16-core CPU (GPU) with similar or lower power consumption. In summary, these results highlight that our FPGA accelerator surpasses general-purpose processors in performance and efficiency, highlighting strong potential for ASIC-based TFHE solutions and practical privacy-preserving computation in real-world applications.